# From Flat to Hierarchical : Extracting Sparse Representations with Matching Pursuit

### v1

@@ Abstract
Motivated by the hypothesis that neural network representations encode abstract, interpretable features as linearly accessible, approximately orthogonal directions, sparse autoencoders (SAEs) have become a popular tool in interpretability literature. However, recent work has demonstrated phenomenology of model representations that lies outside the scope of this hypothesis, showing signatures of hierarchical, nonlinear, and multi-dimensional features. This raises the question: do SAEs represent features that possess structure at odds with their motivating hypothesis? If not, does avoiding this mismatch help identify said features and gain further insights into neural network representations? To answer these questions, we take a construction-based approach and re-contextualize the popular matching pursuits (MP) algorithm from sparse coding to design MP-SAE—an SAE that unrolls its encoder into a sequence of residual-guided steps, allowing it to capture hierarchical and nonlinearly accessible features. Comparing this architecture with existing SAEs on a mixture of synthetic and natural data settings, we show: (i) hierarchical concepts induce conditionally orthogonal features, which existing SAEs are unable to faithfully capture, and (ii) the nonlinear encoding step of MP-SAE recovers highly meaningful features, helping us unravel shared structure in the seemingly dichotomous representation spaces of different modalities in a vision-language model, hence demonstrating the assumption that useful features are solely linearly accessible is insufficient. We also show that the sequential encoder principle of MP-SAE affords an additional benefit of adaptive sparsity at inference time, which may be of independent interest. Overall, we argue our results provide credence to the idea that interpretability should begin with the phenomenology of representations, with methods emerging from assumptions that fit it.
@@ 1 Introduction
Modern neural networks trained on large-scale datasets have achieved unprecedented performance on several practical tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], prompting efforts towards understanding how these abilities are implemented within a model [11, 12, 13]. To this end, Sparse Autoencoders (SAEs) [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], motivated by the Linear Representation Hypothesis (LRH) [24, 25, 25], have become a popular tool for interpreting neural networks. Specifically, LRH, a phenomenological model of organization and computation of neural network representations [24, 25, 25, 26, 27, 28], posits that neural network representations of dimension $m$ can be decomposed along a basis of $p\gg m$ approximately orthogonal directions that reflect abstract, interpretable concepts underlying the data distribution. Here, approximate orthogonality is argued to be necessary for circumventing the problem of packing more unique vectors (concepts) than the dimensionality of the space allows (a.k.a. the superposition problem [24]), hence enabling reliable estimation of a concept’s presence by a linear operator (e.g., the linear map in an MLP) [29, 30]. Grounded in this idea, and inspired by its relation to the well-known problem of sparse coding [31, 32], SAEs have been proposed to learn, in an unsupervised manner, a sparse, overcomplete dictionary of directions that (ideally) map onto abstract, interpretable concepts encoded in a neural network, hence enabling its interpretability [33, 14].
@@
  • Introducing MP-SAE for phenomena beyond linear assumptions. To contextualize the results above and understand if capturing hierarchical structures yields meaningful features in larger-scale models, we design an SAE architecture that can capture (certain kinds of) hierarchical and nonlinear structures. Specifically, we extend the standard SAE framework by unrolling the Matching Pursuit algorithm [74] into a residual-guided, sequential encoder—called the MP-SAE. Each step of MP-SAE selects a feature (approximately) orthogonal to what has already been explained, naturally enforcing a conditionally orthogonal structure. Across experiments on both synthetic benchmarks and real-world vision-language models, MP-SAE proves more expressive and uncovers richer structural features than existing SAEs. These results question the assumption that meaningful features must be linearly and independently accessible.
@@ 2 Formalizing Linear Representation Hypothesis and Sparse Autoencoders
$\begin{cases}\text{{({i})} Overcompleteness:}&p\gg m;\\[2.0pt] \text{{({ii})} Quasi-orthogonality:}&\displaystyle\max_{i\neq j}\bigl{|}\bm{D}_{i}^{\!\top}\bm{D}_{j}\bigr{|}\leq\varepsilon,\quad\text{where }\forall i~{}~{}\|\bm{D}_{i}\|_{2}=1\ ;\text{ and}\\ \text{{({iii})} Sparsity:}&|\operatorname{supp}(\bm{z})|\leq k\ll p.\\ \end{cases}$
@@ 2.1 Stress-Testing SAEs via Conditional Orthogonality—a Structure Outside LRH’s Scope
Figure 2: Illustrative Example of Conditional vs. Quasi-Orthogonality. A) Example of a hierarchical concept tree. B) Comparison of quasi-orthogonality (interference within levels) vs. conditional orthogonality (orthogonality across levels). C) Correlation matrix of features sampled from A, showing conditional orthogonality (white, $=0$) across levels and quasi-orthogonality (light blue, $=\varepsilon$) within levels.
@@ 3 Operationalizing Conditional Orthogonality via MP-SAE
To enable the ability to identification of features encoded in a conditionally orthogonal manner (Def. 2.3), we construct a sparse autoencoder whose inference reflects the structure of hierarchical representations—minimizing interference across levels while tolerating it within [69]. Our design draws on the Matching Pursuit (MP) algorithm [74] from sparse coding [89, 90, 91, 92] and hierarchical Bayesian modeling [93]. In these settings, inference proceeds greedily: at each step, the feature most correlated with the residual is selected, its contribution subtracted, and the process repeated. This iterative residual decomposition promotes conditional orthogonality, as each new feature explains variance orthogonal to what has already been captured (Fig. 3). Embedding this mechanism into a sparse autoencoder yields the Matching Pursuit Sparse Autoencoder (MP-SAE), formalized in Algorithm 1. Below, we unpack the structure and implications of MP-SAE, examining: (i) the mechanics of its inference procedure; (ii) how this procedure promotes conditional orthogonality; and (iii) its capacity to extract higher-order, nonlinearly accessible features.
@@
Figure 3: (Left) One iteration of the MP-SAE encoder. At each iteration, the residual $\bm{r}^{(t)}$ is projected onto the dictionary $\bm{D}$, and the most correlated feature $\bm{D}_{j^{(t)}}$ is selected. Its contribution is subtracted from the residual and added to the reconstruction. (Right) MP-SAE Algorithm. At each iteration, a concept is selected to explain the residual $\bm{r}$—the part of $\bm{x}$ not yet reconstructed—yielding a sequential decomposition. This residual-guided, greedy inference promotes conditional orthogonality and enables access to higher-order features that are nonlinearly accessible from $\bm{x}$.
@@
MP-SAE instantiates the classical Matching Pursuit algorithm [74] as a neural encoder by unrolling its greedy inference procedure into a fixed number of steps. At inference, the model starts with an initial estimate and residual (Algorithm 1, lines 3 and 2). At each iteration, MP-SAE greedily selects the feature that best aligns with the current residual by computing the inner product between the residual and each feature, and choosing the one with the highest projection (line 6). Once the best-matching feature is identified, the algorithm determines its contribution by projecting the residual onto that feature (line 7), adds this contribution to the current approximation of the input (line 8), and updates the residual accordingly (line 9). This procedure is repeated for $T$ steps, producing a sparse code with $\|\bm{z}\|_{0}\leq T$. The resulting encoding represents a sequential approximation of the input, constructed through greedy selection of locally optimal features that progressively refine the reconstruction of $\bm{x}$.
@@
A defining property of MP-SAE is that its greedy inference procedure promotes conditional orthogonality across selected features. This emerges from the residual update rule: once a concept $\bm{D}_{j^{(t-1)}}$ has been selected from the dictionary at iteration $t{-}1$, we find that the updated residual ${\bm{r}}^{(t)}$ is orthogonal to it, as stated below formally.
@@
Let ${\bm{r}}^{(t)}$ be the residual at iteration $t$ of MP-SAE inference, and let $\bm{D}_{j^{(t-1)}}$ be the feature selected at step $t{-}1$. Then:
@@
Each selected concept is removed from the residual subspace before the next selection, and subsequent concepts are chosen to explain what remains in the residual, orthogonal to the last selected feature. When trained, this yields a sequence of features that are not globally orthogonal, but conditionally orthogonal. Although MP-SAE only enforces orthogonality to the most recently selected concept—unlike Orthogonal Matching Pursuit [89], which re-orthogonalizes against all prior selections—we observe empirically that residuals are often nearly orthogonal to all previously selected directions. This emergent behavior suggests the model implicitly promotes hierarchical separation in the learned dictionary, minimizing interference across levels. It aligns with the inductive bias of conditional orthogonality (Def. 2.3) and contrasts with standard SAEs, which enforce global quasi-orthogonality regardless of feature structure.
(iii) Access higher-order nonlinear features
$\bm{x}={\bm{r}}^{(0)}=\bm{\varphi}(\bm{x})+{\bm{r}}^{(1)}=\underbrace{\bm{\varphi}(\bm{x})}_{\text{{linearly accessible}}}+\underbrace{\sum_{t=1}^{T}\bm{\varphi}({\bm{r}}^{(t)})}_{\text{{nonlinearly accessible}}}+\underbrace{{\bm{r}}^{(T+1)}}_{\text{{residual error}}},\vspace{-2mm}$
where $\bm{\varphi}(\cdot)$ denotes the linear projection onto the selected feature at each step (See Figure 3). The crucial insight is that, although each $\bm{\varphi}({\bm{r}}^{(t)})$ is linear in its argument, the argument itself, ${\bm{r}}^{(t)}$, is a nonlinear function of $\bm{x}$. As a result, the composition $\bm{\varphi}({\bm{r}}^{(t)})$ defines a feature that cannot be obtained from $\bm{x}$ via a single linear map. MP-SAE thus naturally enables the discovery of higher-order concepts that are conditionally dependent on previously explained structure. This mechanism may be particularly relevant in settings where important features are entangled or nonlinearly composed such as hierarchies, temporal dependencies, or multimodal correlations. Moreover, it provides a constructive hypothesis for the phenomenon of “dark matter” in neural representations [71]: features that evade standard SAEs because they are not linearly accessible from the raw representation; that is, one can still have the LRH assumption of a linear mixing process hold, i.e., $x=Dz$, but other constraints start to relax (Def. 2.1). We return to this phenomenon in our empirical analysis.
4 Results
We now take a step towards uncovering challenges in SAEs emergent from assuming only a partially correct model of neural network representations, i.e., LRH. Specifically, in Sec. 4.1, we analyze a synthetic domain that helps highlight the inability of existing SAEs to identify hierarchically structured concepts in a clearly defined domain. Building on these results, in Sec. 4.2, we analyze a natural vision-language domain that helps us assess how features extracted under an inductive bias that accommodates hierarchical, nonlinear concepts—as done in MP-SAE—are differently from ones identified using existing SAEs.
4.1 A Synthetic Generative Model of Hierarchical Features
@@
We first evaluate whether SAEs (including our proposed MP-SAE) can recover conditionally orthogonal features from a synthetic hierarchy-based benchmark adapted from Bussmann et al. [20]. Specifically, we analyze Vanilla (ReLU) [23, 14], BatchTopK [17], and Matryoshka [20, 94] all trained with fixed $\ell_{0}$ sparsity targets. The ground truth consists of 20 unit-norm concepts organized into a two-level tree: 11 disjoint parent concepts, 3 of which each have 3 children. In the low-interference regime (Fig. 4A), Vanilla and BatchTopK suffer from feature absorption [95], aligning child concepts with their parent and collapsing hierarchy—though they retain within-level corelation. Matryoshka avoids absorption and preserves hierarchy but introduces negative interference between siblings, distorting flat structure. Only MP-SAE recovers both intra and inter level structure. To further stress-test the different SAEs, we vary within-level correlation in the ground truth and evaluate learned dictionaries using “Flat MSE” (for intra-level alignment; see Fig. 4B) and “Hierarchical MSE” (for inter-level separation; see Fig. 4C). We observe that Matryoshka tends to preserves hierarchy but loses flat structure; meanwhile, Vanilla and BatchTopK behave oppositely. MP-SAE exhibits three regimes: low to moderate interference yields recovery of both structures; under high interference, both degrade, reaching eventually a point where we see MP-SAE sacrifice flatness to maintain hierarchy, highlighting its inductive bias.
@@
Overall, this synthetic benchmark elicits challenges towards capturing hierarchical features via existing SAEs, including ones trained with explicit hierarchy-promoting objectives. Instead, we see capturing the relevant inductive bias via MP-SAE works really well: MP-SAE is uniquely capable of recovering hierarchically structured features under controlled conditions. We now turn to pretrained representations to assess whether such structure arises in practice and how SAEs respond to it.

### v2

@@ Abstract
Motivated by the hypothesis that neural network representations encode abstract, interpretable features as linearly accessible, approximately orthogonal directions, sparse autoencoders (SAEs) have become a popular tool in interpretability literature. However, recent work has demonstrated phenomenology of model representations that lies outside the scope of this hypothesis, showing signatures of hierarchical, nonlinear, and multi-dimensional features. This raises the question: do SAEs represent features that possess structure at odds with their motivating hypothesis? If not, does avoiding this mismatch help identify said features and gain further insights into neural network representations? To answer these questions, we take a construction-based approach and re-contextualize the popular matching pursuit (MP) algorithm from sparse coding to design MP-SAE—an SAE that unrolls its encoder into a sequence of residual-guided steps, allowing it to capture hierarchical and nonlinearly accessible features. Comparing this architecture with existing SAEs on a mixture of synthetic and natural data settings, we show: (i) hierarchical concepts induce conditionally orthogonal features, which existing SAEs are unable to faithfully capture, and (ii) the nonlinear encoding step of MP-SAE recovers highly meaningful features, helping us unravel shared structure in the seemingly dichotomous representation spaces of different modalities in a vision-language model, hence demonstrating the assumption that useful features are solely linearly accessible is insufficient. We also show that the sequential encoder principle of MP-SAE affords an additional benefit of adaptive sparsity at inference time, which may be of independent interest. Overall, we argue our results provide credence to the idea that interpretability should begin with the phenomenology of representations, with methods emerging from assumptions that fit it.
@@ 1 Introduction
Modern neural networks trained on large-scale datasets have achieved unprecedented performance on several practical tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], prompting efforts towards understanding how these abilities are implemented within a model [11, 12, 13]. To this end, Sparse Autoencoders (SAEs) [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], motivated by the Linear Representation Hypothesis (LRH) [24, 25], have become a popular tool for interpreting neural networks. Specifically, LRH, a phenomenological model of organization and computation of neural network representations [24, 25, 26, 27, 28], posits that neural network representations of dimension $m$ can be decomposed along a basis of $p\gg m$ approximately orthogonal directions that reflect abstract, interpretable concepts underlying the data distribution. Here, approximate orthogonality is argued to be necessary for circumventing the problem of packing more unique vectors (concepts) than the dimensionality of the space allows (a.k.a. the superposition problem [24]), hence enabling reliable estimation of a concept’s presence by a linear operator (e.g., the linear map in an MLP) [29, 30]. Grounded in this idea, and inspired by its relation to the well-known problem of sparse coding [31, 32], SAEs have been proposed to learn, in an unsupervised manner, a sparse, overcomplete dictionary of directions that (ideally) map onto abstract, interpretable concepts encoded in a neural network, hence enabling its interpretability [33, 14].
@@
  • Introducing MP-SAE for phenomena beyond linear assumptions. To contextualize the results above and understand if capturing hierarchical structures yields meaningful features in larger-scale models, we design an SAE architecture that can capture (certain kinds of) hierarchical and nonlinear structures. Specifically, we extend the standard SAE framework by unrolling the Matching Pursuit algorithm [74] into a residual-guided, sequential encoder—called the MP-SAE. Each step of MP-SAE selects a feature (approximately) orthogonal to what has already been explained, naturally promoting a conditionally orthogonal structure. Across experiments on both synthetic benchmarks and real-world vision-language models, MP-SAE proves more expressive and uncovers richer structural features than existing SAEs. These results question the assumption that meaningful features must be linearly and independently accessible.
@@ 2 Formalizing Linear Representation Hypothesis and Sparse Autoencoders
[…]
@@ 2.1 Stress-Testing SAEs via Conditional Orthogonality—a Structure Outside LRH’s Scope
Figure 2: Illustrative Example of Conditional vs. Quasi-Orthogonality. A) Example of a hierarchical concept tree. B) Comparison of quasi-orthogonality (interference within levels) vs. conditional orthogonality (orthogonality across levels). C) Correlation matrix of features sampled from A, showing conditional orthogonality (white, =0=0) across levels and quasi-orthogonality (light blue, $\varepsilon$) within levels.
@@ 3 Operationalizing Conditional Orthogonality via MP-SAE
To enable the ability to identification of features encoded in a conditionally orthogonal manner (Def. 2.3), we construct a sparse autoencoder whose inference reflects the structure of hierarchical representations—minimizing interference across levels while tolerating it within [69]. Our design draws on the Matching Pursuit (MP) algorithm [74] from sparse coding [89, 90, 91, 92], which has been shown to recover features from the appropriate hierarchical level in conditionally orthogonal dictionaries [93]. MP performs inference greedily: at each step, the feature most correlated with the residual is selected, its contribution subtracted, and the process repeated. This iterative residual decomposition promotes conditionally orthogonal feature selection, as each new feature explains variance orthogonal to what was captured by the previous feature (Fig. 3). Embedding this mechanism into a sparse autoencoder yields the Matching Pursuit Sparse Autoencoder (MP-SAE), formalized in Algorithm 1. Below, we unpack the structure and implications of MP-SAE, examining: (i) the mechanics of its inference procedure; (ii) how this procedure promotes conditional orthogonality; and (iii) its capacity to extract higher-order, nonlinearly accessible features.
@@
Figure 3: (Top Left) One iteration of the MP-SAE encoder: The residual $\bm{r}^{(t)}$ is projected onto the dictionary $\bm{D}$, the most correlated feature $\bm{D}_{j^{(t)}}$ is selected, and its contribution is subtracted from the residual and added to the reconstruction. (Bottom Left) MP sequentially reconstructs $\bm{x}$ by greedily selecting features that best explain the residual, promoting orthogonal selection and enabling access to higher-order features that are nonlinearly accessible from $\bm{x}$. (Right) MP-SAE embeds the MP algorithm into a sparse autoencoder, where the dictionary $\bm{D}$ is learned through backpropagation.
@@
MP-SAE uses a shared learned dictionary for encoding and decoding. The encoder embeds the classical Matching Pursuit algorithm [74] by unrolling its greedy inference procedure into a fixed number of steps. At inference, the model starts with an initial estimate and residual (Algorithm 1, lines 3 and 2). At each iteration, MP-SAE greedily selects the feature that best aligns with the current residual by computing the inner product between the residual and each feature, and choosing the one with the highest projection (line 6). Once the best-matching feature is identified, the algorithm determines its contribution by projecting the residual onto that feature (line 7), adds this contribution to the current approximation of the input (line 8), and updates the residual accordingly (line 9). This procedure is repeated for $T$ steps, producing a sparse code with $\|\bm{z}\|_{0}\leq T$. The resulting encoding represents a sequential approximation of the input, constructed through greedy selection of locally optimal features that progressively refine the reconstruction of $\bm{x}$.
@@
A defining property of MP-SAE is that its greedy inference procedure promotes conditional orthogonality across selected features. This emerges from the residual update rule: once a concept $\bm{D}_{j^{(t-1)}}$ has been selected from the dictionary at iteration t1t{-}1, the updated residual ${\bm{r}}^{(t)}$ is orthogonal to it, as stated below formally.
@@
Let ${\bm{r}}^{(t)}$ be the residual at iteration $t$ of MP-SAE inference, and let $\bm{D}_{j^{(t-1)}}$ be the feature selected at step t1t{-}1. Then:
@@
Each selected concept is removed from the residual subspace before the next selection, and subsequent concepts are chosen to explain what remains in the residual, orthogonal to the last selected feature. When trained, this yields a dictionary of features that are not globally orthogonal but are selectively chosen to be conditionally orthogonal. Although MP-SAE only enforces orthogonality to the most recently selected concept—unlike Orthogonal Matching Pursuit [89], which re-orthogonalizes against all prior selections—we observe empirically that residuals are often nearly orthogonal to all previously selected directions. This emergent behavior suggests the model implicitly promotes hierarchical separation in the learned dictionary, minimizing interference across levels. It aligns with the inductive bias of conditional orthogonality (Def. 2.3) and contrasts with standard SAEs, which tend to promote global quasi-orthogonality regardless of feature structure.
(iii) Access nonlinearly accessible features.
$\bm{x}={\bm{r}}^{(0)}=\bm{\varphi}(\bm{x})+{\bm{r}}^{(1)}=\underbrace{\bm{\varphi}(\bm{x})}_{\text{{linearly accessible}}}+\underbrace{\sum_{t=1}^{T}\bm{\varphi}({\bm{r}}^{(t)})}_{\text{{nonlinearly accessible}}}+\underbrace{{\bm{r}}^{(T+1)}}_{\text{{residual error}}},\vskip-5.69054pt$
where $\bm{\varphi}(\cdot)$ denotes the linear projection onto the selected feature at each step (See Fig. 3). The crucial insight is that, although each $\bm{\varphi}({\bm{r}}^{(t)})$ is linear in its argument, the argument itself, ${\bm{r}}^{(t)}$, is a nonlinear function of $\bm{x}$. As a result, the composition $\bm{\varphi}({\bm{r}}^{(t)})$ defines a feature that cannot be obtained from $\bm{x}$ via a single linear map. MP-SAE can thus potentially uncover higher-order concepts that are conditionally dependent on previously explained structure. This mechanism may be particularly relevant in settings where important features are entangled or nonlinearly composed such as hierarchies, temporal dependencies, or multimodal correlations. Moreover, it provides a constructive hypothesis for the phenomenon of “dark matter” in neural representations [71]: features that evade standard SAEs because they are not linearly accessible from the raw representation; that is, one can still have the LRH assumption of a linear mixing process hold, i.e., $x=Dz$, but other constraints start to relax (Def. 2.1). We return to this phenomenon in our empirical analysis.
4 Empirical Results
We now take a step towards uncovering challenges in SAEs emergent from assuming a partially correct model of neural network representations, i.e., LRH. Specifically, in Sec. 4.1, we analyze a synthetic domain that highlight the inability of existing SAEs to identify hierarchically structured concepts in a clearly defined domain. Building on these results, in Sec. 4.2, we analyze a natural vision–language domain to assess how features extracted under an inductive bias that accommodates hierarchical, nonlinearly accessible concepts—as in MP-SAE—differ from those identified by existing SAEs. In particular, we show that MP-SAE uniquely recovers cross-modal structure.
4.1 A Synthetic Generative Model of Hierarchical Features
We begin by evaluating whether SAEs, including our proposed MP-SAE, can recover conditionally orthogonal features using a synthetic hierarchy-based setup adapted from Bussmann et al. [20]. To formalize this setting, we introduce the following definition of a hierarchical generative process.
Definition 4.1 (Hierarchical Generative Process).
A generative process over $\bm{p}$ (parent concept) and $\bm{c}$ (child concept) is said to be hierarchical if their activations $z_{p}$ and $z_{c}$ satisfy
$\mathbb{P}(z_{p}>0\mid z_{c}>0)=1\quad\text{and}\quad\mathbb{P}(z_{c}>0\mid z_{p}>0)<1.$
That is, a child’s activation requires its parent’s, whereas a parent may activate independently.
Based on this definition, we implemented the following hierarchical generative process. The ground truth $\bm{D}$ consists of 20 unit-norm concepts organized into a two-level tree: 11 disjoint parent concepts $\bm{D}_{p}$, 3 of which each have 3 disjoint children $\bm{D}_{c}$ as depicted in Fig. 2. Each input is generated by sampling a parent and, if present, a child, resulting in one or two active concepts.
$\bm{x}=z_{p}\cdot\bm{D}_{p}+z_{c}\cdot\bm{D}_{c}$
The activations $z_{p}$ and $z_{c}$ follow the activation pattern defined in Definition 4.1, taking values from $\mathcal{N}(1.5,\nicefrac{{1}}{{4^{2}}})$ (ensuring positive values) when active and 0 otherwise. [^2] We generalize the generative process from Bussmann et al. [20], where $z_{p}$ and $z_{c}$ are fixed across all inputs, making their firing magnitudes perfectly correlated ($z_{p}=\lambda z_{c}$). See Appendix B.1.4 for an extended discussion.
@@ 4.1 A Synthetic Generative Model of Hierarchical Features
We analyze Vanilla (ReLU) [23, 14], BatchTopK [17], and Matryoshka [20, 94] all trained with fixed $\ell_{0}$ sparsity targets. In a fixed low intra-level interference regime (Fig. 4A), Vanilla and BatchTopK suffer from feature absorption [95], aligning child concepts with their parent and collapsing hierarchy—though they retain within-level corelation. Matryoshka avoids absorption and preserves hierarchy but introduces negative interference between siblings, distorting flat structure. Only MP-SAE recovers both intra and inter level structure. To further stress-test the different SAEs, we vary within-level correlation in the ground truth and evaluate learned dictionaries using “Flat MSE” (for intra-level alignment; see Fig. 4B) and “Hierarchical MSE” (for inter-level separation; see Fig. 4C). We observe that Matryoshka tends to preserves hierarchy but loses flat structure; meanwhile, Vanilla and BatchTopK behave oppositely. MP-SAE exhibits three regimes: low to moderate interference yields recovery of both structures; under high interference, both degrade, reaching eventually a point where we see MP-SAE sacrifice flatness to maintain hierarchy, highlighting its inductive bias.
@@
Overall, this synthetic benchmark elicits challenges towards capturing hierarchical features via existing SAEs, including ones trained with explicit hierarchy-promoting objectives. Instead, capturing the relevant inductive bias via MP-SAE works really well: MP-SAE is uniquely capable of recovering hierarchically structured features when such property statistically exists. We now turn to pretrained representations to assess whether such structure arises in practice and how SAEs respond to it.